OpenStudy (anonymous):

Use the Change of Base Formula to evaluate log5 44. Then convert log5 44 to a logarithm in base 2. Round to the nearest thousandth.

OpenStudy (anonymous):

\[\log_5(44)=\frac{\log(44)}{\log(5)}\]

OpenStudy (anonymous):

it is always true that \[\log_b(x)=\frac{\log_a(x)}{\log_a(b)}\] but if you want to get a decimal with a calculator, you have to use \(\log_{10}(x)\) or \(\ln(x)\)

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."